OpenStudy (anonymous):

Find the values of x in this equation: x-15/x=2 A. -7, 3 B. -5, 2 C. -7, 5 D. -2, 5 E. -3, 5

OpenStudy (anonymous):

x^2 - 15 = 2x x^2 - 2x - 15 = 0 (x-5)*(x+3) x is -3 and 5 E

OpenStudy (anonymous):

THank you so much im in desperate need of help im taking an online exam for pre calc and I need to pass it for my diplomia:(

OpenStudy (anonymous):

i do not condone cheating

OpenStudy (anonymous):

its not cheating if my principle gives me permission to find help.

OpenStudy (anonymous):

then what's the point of giving you a test if you're not the one answering?

OpenStudy (anonymous):

I failed my pre calc exam during HS iwasnt able to walk the stage so he gave me an online class to take so I could get my diploma over the summer.

OpenStudy (anonymous):

he told me I can get help if needed. as long as he sees me get my diploma hes happy.

OpenStudy (anonymous):

its not right for you to call me out for "cheating" when 90% of people on here are getting help posting there questions to there tests or homework.

OpenStudy (anonymous):

again thank you for your help good day.

OpenStudy (anonymous):

thank you for your rant. There is a difference on getting help on homework and getting help on a test that depends on whether you graduation high school or not. I hope you someday you will learn the value of hard work and the reward of passing high school based on your own knowledge, not somebody else's